For two consecutive integers the sum of the smaller and twice the larger is 38 what are the integers?
Answer by mananth(16949)

let the integers be x & x+1
x+2(x+1)=38
x+2x+2=38
3x+2=38
3x=38-2
3x=36
/3
x=12
The numbers are 12, 13